OpenStudy (anonymous):

What is the sum of the first 30 terms of this arithmetic sequence? 6, 13, 20, 27, 34, … PLEASE HELP

OpenStudy (solomonzelman):

\(\large\color{slate}{ d=(a_n)-(a_{n-1}) }\) find the difference:

OpenStudy (anonymous):

A. 3225 B.3255 C. 3325 D.3525

OpenStudy (solomonzelman):

Find the difference (d), with the formula I gave; \(\large\color{slate}{ a_{30}=a_1+d(30-1)=? }\) (do this after knwoing the d) then use: \(\large\color{slate}{ S_{30}=(1/2)\cdot(30)\cdot(a_1+a_{30}) }\)

OpenStudy (anonymous):

I got b, this is a pretest

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

is that right

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

nm its a

OpenStudy (solomonzelman):

so a or b?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

yes it is A.

OpenStudy (solomonzelman):

I was absent sorry for replying late